# Cellular, Cell-less, and Everything in Between: A Unified Framework for Utility Region Analysis in Wireless Networks

### v2

@@ IV Discussion
To highlight further the relevance of the preceding results for network designers, except for Sect. IV-E, we focus on the interference model in Examples 3 and 4. Notably, that interference model is very common in the massive/extremely large MIMO and cell-less literature (see [4, 25, 5] and the references therein). In the discussion that follows, we extensively use [30, Theorem 2.1] (Fact 4 in the Supplemental Material), which establishes that an invertible interference matrix ${\bm{M}}\in{\mathbb{R}}_{+}^{N\times N}$, being nonnegative, is an inverse Z-matrix if and only it is an inverse M-matrix. This relationship allows us to utilize the insights from the previous sections, along with the comprehensive body of research on these matrix classes [30], to explain interference patterns observed in common network models for resource allocation.
@@ IV-A Impact of self-interference on the convexity of the feasible rate and SINR regions
The diagonal entries of the interference matrix ${\bm{M}}\in{\mathbb{R}}_{+}^{N\times N}$ defined in Example 4 corresponds to the so-called self-interference, which in many models in the massive MIMO and cell-less literature is caused by the lack of perfect channel state information. For Remark 2 to be applicable, self-interference must exist in all communication links $n\in\mathcal{N}$. The reason is that a necessary (but not sufficient) condition for ${\bm{M}}\in{\mathbb{R}}_{+}^{N\times N}$ to be an inverse Z-matrix (or, equivalently, an inverse M-matrix, as mentioned above) is that all diagonal entries must be positive [30, Theorem 2.9].
@@
The results in [30, Theorems 4.5 and 9.2] (Fact 2 in the Supplemental Material) show that there exists scalars $(\alpha_{1},\ldots,\alpha_{K+1})\in{\mathbb{R}}_{+}^{K+1}$ such that $(\alpha_{n}{\bm{I}}+{\bm{M}}+{\bm{u}}{\bm{a}}_{n}^{t})_{n\in\{1,\ldots,K\}}$ and $\alpha_{K+1}{\bm{I}}+{\bm{M}}$ are inverse M-matrices. It then follows from [30, Theorem 1.7] (Fact 3 in the Supplemental Material) that, for any ${\bm{S}}\in\mathcal{D}_{+}^{N\times N}$ satisfying ${\bm{S}}\geq{\bm{S}}_{\mathrm{min}}:=\bar{\alpha}{\bm{I}}$, where $\bar{\alpha}:=\max\limits_{n\in\{1,\ldots,K+1\}}\alpha_{n}$, the matrices ${\bm{M}}+{\bm{S}}$, ${\bm{M}}+{\bm{S}}+{\bm{u}}{\bm{a}}_{1}^{t}$, …, ${\bm{M}}+{\bm{S}}+{\bm{u}}{\bm{a}}_{K}^{t}$ are inverse M-matrices, hence also inverse Z-matrices because they are nonnegative [30, Theorem 2.1] (Fact 4 in the Supplemental Material). A direct application of Remark 2, Corollary 1, and Corollary 2 completes the proof. ∎
@@ IV-B Users offering compatible channels
Assume that two single-antenna users need to be served by a single basestation equipped with $L\in{\mathbb{N}}$ antennas (we consider the general case soon). In the massive MIMO literature, a key concept indicating whether both users can be served simultaneously while maintaining good link performance is favorable or approximately favorable propagation [7, Ch. 7]. More precisely, we say that the users offer favorable propagation if their channels $({\bm{h}}_{1},{\bm{h}}_{2})\in\mathbb{C}^{L}\times\mathbb{C}^{L}$ are such that ${\bm{h}}_{1}^{H}{\bm{h}}_{2}=0$, or that they offer approximately favorable propagation if $|{\bm{h}}_{1}^{H}{\bm{h}}_{2}|$ is sufficiently small (possibly in a probabilistic sense for models such as those in Example 3). It is often recommended to avoid scheduling users on the same resources when favorable propagation is absent [7, p. 154], and this principle can be trivially generalized to systems with multiple users and basestations by considering the channels for each link pair. However, this system design recommendation overlooks key aspects of real systems, including, among others: (i) it does not take it account the presence of beamformers and the available transmit power; (ii) it ignores self-interference; (iii) many results proving approximate favorable propagation consider asymptotic regimes by letting the number of antennas diverge to infinity, but systems with a small number of antennas are common; (iii) users cannot be simply dropped – they need to be scheduled at some time slot considering some fairness criterion; (iv) and it lacks mathematical guarantees that users would not achieve better performance by being divided into smaller groups and served at different time slots.
The results in Sect. III provide a framework for defining a notion of user compatibility that overcomes the above drawbacks. Specifically, if the achievable rate region is convex, as discussed in Sect. I-A, we have a mathematical certificate that the network performance cannot be improved by splitting users into smaller groups and by serving these groups at different times, irrespective of the fairness criterion being considered. Therefore, the following definition is justified:
@@
Consider the system model in Examples 3 and 4. We say that the users offer Z-compatible channels without power constraints if the interference matrix $\bm{M}$} is an inverse Z-matrix. Likewise, we say that the users offer Z-compatible channels with the power constraint $\mathcal{P}$ in (18) if the matrices $({\bm{M}}+{\bm{ua}}_{n}^{t})_{n\in\{1,\ldots,K\}}$ are inverse Z-matrices.
@@ IV-C Numerical aspects
If the interference matrix $\bm{M}$} is of dimension $2\times 2$ (i.e., we try to produce a certificate of channel Z-compatibility for two users), inverting such a small matrix may not pose serious challenges. Nevertheless, it is more efficient and numerically stable to apply the result stated in [30, Theorem 2.9.2] (Fact 1 in the Supplemental Material). More specifically, an interference matrix ${\bm{M}}\in{\mathbb{R}}_{+}^{2\times 2}$ is an inverse Z-matrix if and only if its determinant $\mathrm{det}({\bm{M}})$ satisfies $\mathrm{det}({\bm{M}})>0$. Unfortunately, this simple inequality is only a necessary condition for larger matrices [30, Theorem 1.2]. Nevertheless, all principal minors of inverse Z-matrices are positive [30, Corollary 2.3.3], which in practice means that a necessary (but not sufficient) condition for Z-compatibility of all channels within a group of users is that any two channels are Z-compatible, which lead us to a simple rule for identifying users within a group potentially leading to a nonconvex SINR or rate region:
@@
Consider the interference model in Examples 3 and 4. Let ${\bm{M}}\in{\mathbb{R}}_{+}^{N\times N}$ be the interference matrix of a network with NN users. Given two users $(i,j)\in\mathcal{N}\times\mathcal{N}$, let ${\bm{M}}[i,j]\in{\mathbb{R}}^{2\times 2}_{+}$ be the reduced interference matrix obtained from $\bm{M}$} by keeping only the $i$th and $j$th columns and the $i$th and $j$th rows. If $\mathrm{det}({\bm{M}}[i,j])\leq 0$, then the full matrix $\bm{M}$} is not an inverse Z-matrix.
IV-D Disproving the conjecture in [13, Sect. 5.4.4][24]
Given a matrix ${\bm{M}}\in{\mathbb{R}}_{+}^{N\times N}$, the studies in [13, Sect. 5.4.4][24] conjecture that convexity of the function $l_{\bm{M}}:{\mathbb{R}}^{N}_{+}\to{\mathbb{R}}_{+}:{\bm{x}}\mapsto\rho({\mathrm{diag}(\bm{{\bm{x}}})}{\bm{M}})$ (and, hence, convexity of the SINR region $\mathcal{S}$ for the system model in Example 3) is linked to the positive semidefiniteness of the symmetrized interference matrix ${\bm{M}}+{\bm{M}}^{t}$. We now show that this condition is neither necessary nor sufficient. More specifically, if the interference matrix ${\bm{M}}\in{\mathbb{R}}^{N\times N}_{+}$ is an inverse Z-matrix, in which case the function $l_{\bm{M}}$ is convex, then there exists a positive diagonal matrix ${\bm{D}}\in{\mathbb{R}}^{N\times N}_{+}$ such that the symmetrized interference matrix ${\bm{DM}}+{\bm{M}}^{t}{\bm{D}}$ is positive definite [30, Theorem 1.4], but $\bm{D}$} is not necessarily the identity matrix. Therefore, positive definiteness of ${\bm{M}}+{\bm{M}}^{t}$ is not a necessary condition for convexity of $l_{\bm{M}}$. To show that this property is not sufficient, take ${\bm{x}}_{1}=[0.5,~0.1,~1]^{t}$, ${\bm{x}}_{2}=[0.5,~0.5,~0.5]^{t}$, $\alpha=0.9$, and
@@ IV-E The sum-rate maximization problem
In this final subsection of Sect. IV, we turn the attention to the interference model in (15) (or, equivalently (17)) in its full generality, with the models presented in Examples 3 and 4 serving as particular cases. In more detail, many scheduling and resource allocation schemes rely heavily on the sum-rate maximization problem, which is known to be NP-hard in general [14]. However, if $T_{\|\cdot\|}$} in (20) is norm inducing, the closure $\overline{\mathcal{R}_{\mathcal{P}}}=\{{\bm{r}}\in{\mathbb{R}}_{+}^{N}\mid\rho((e^{{\mathrm{diag}(\bm{r})}}-{\bm{I}}){T_{\|\cdot\|}})\leq 1\}$ of the set of achievable rates $\mathcal{R}_{\mathcal{P}}$ in (27) is convex as shown in Proposition 6, and $\overline{\mathcal{R}_{\mathcal{P}}}$ can be equivalently expressed using a monotone norm as shown in (29). In this case, (weighted) sum-rate maximization can be posed as the following convex optimization problem, which opens up the possibility for deriving efficient solvers that provably converge to global optima (particular solvers will be considered elsewhere):
$\begin{array}[]{rl}\text{maximize}&{\bm{w}}^{t}{\bm{r}}\\ \text{subject~to}&{\bm{r}}\in\overline{\mathcal{R}_{\mathcal{P}}}\end{array}$ (32)
@@
It is worth emphasizing that the class of sum-rate maximization problems that can be reformulated as convex optimization problems does not necessarily require $T_{\|\cdot\|}$} to be norm inducing, and the results from Sect. III offer analytical support for a straightforward guideline for designing sum-rate maximization algorithms:
@@
In Proposition 6, we have shown that convexity of the function ${\bm{s}}\mapsto\rho({\mathrm{diag}(\bm{s})}{T_{\|\cdot\|}})$ implies convexity of the sets $\overline{{\mathcal{S}}_{\mathcal{P}}}$ and $\overline{\mathcal{R}_{\mathcal{P}}}$. However, the arguments cannot be reversed in general: $\overline{\mathcal{R}_{\mathcal{P}}}$ can be convex even if neither the set $\overline{{\mathcal{S}}_{\mathcal{P}}}$ nor the function ${\bm{s}}\mapsto\rho({\mathrm{diag}(\bm{s})}{T_{\|\cdot\|}})$ is convex. This fact suggests to use the achievable rates directly as the optimization variables in sum-rate maximization problems, as done in (32). The alternative approach of using SINR values ${\bm{s}}=(s_{1},\ldots,s_{N})$ as the optimization variables (or the power levels providing the required SINR values) and the sum-rate function $\sum_{n\in\mathcal{N}}\log(1+s_{n})$ as the cost function may require the introduction of the set $\overline{{\mathcal{S}}_{\mathcal{P}}}$ as a constraint, which, for the reasons mentioned above, is less likely to be convex compared to the set $\overline{\mathcal{R}_{\mathcal{P}}}$. Therefore, we may lose the opportunity to exploit a potential hidden convexity of the sum-rate maximization problem, increasing the risk of iterative solvers converging to local rather than global optima, if convergence can indeed be guaranteed.
@@ V Simulations
To illustrate the results from the previous section using the model in Examples 3 and 4, we examine the uplink of a small cell-less network consisting of four access points, each equipped with two antennas, uniformly distributed within a 100m × 100m square area. Three single-antenna users are positioned uniformly at random in this area, with each user connected to the two access points providing the strongest channels. The small number of users and access points is chosen deliberately, because plotting the feasible SINR and rate regions is impossible if the number of users exceeds three. Nevertheless, the results in the previous section are valid for systems of any dimension. For the beamformers, we employ maximum ratio combining. Concretely, in the SINR expression in (16), the aggregate beamformer vector ${\bm{v}}_{n}$ (obtained by stacking the beamformers of every access point) for each user $n\in\mathcal{N}=\{1,2,3\}$ is a scaled version of the aggregated channel vector ${\bm{h}}_{n}$, with zeros placed in the coordinates corresponding to the access points to which the user is not connected. Expectations are replaced by empirical averages from 100 realizations of the random variables, and all channel parameters and noise samples are constructed exactly as described in [4, Sect. III-D], so we omit the details for brevity. For the monotone norm $\|\cdot\|$ in the power constraint $\mathcal{P}$ in (18), we use the $l_{\infty}$ norm, scaled such that the maximum transmit power per user is limited to $p_{\max}=0.2$ W; i.e., we use the polyhedral monotone norm $(\forall{\bm{p}}\in{\mathbb{R}}^{N})~\|{\bm{p}}\|:=(1/p_{\max})~\|{\bm{p}}\|_{\infty}$. Equivalently, the vectors $({\bm{a}}_{n})_{n\in\{1,2,3\}}$ in Corollary 2 are the standard basis vectors appropriately scaled so that the users do not exceed the maximum transmit power.
@@
Fig. 2 illustrates results for a sample user placement scenario where all matrices $({\bm{M}}+{\bm{u}}{\bm{a}}^{t}_{n})_{n\in\mathcal{N}}$ described in Example 3 are inverse Z-matrices. These figures indicate convexity of the feasible rate and SINR regions, which is in agreement with the results in Corollary 2.
@@
which, in light of Remark 4, cannot be an inverse Z-matrix because, in particular, the determinant of the $2\times 2$ matrix obtained by removing the second row and the second column is negative; i.e., users 1 and 3 do not offer Z-compatible channels. Consequently, our theoretical results cannot guarantee convexity of the SINR and rate regions, and we indeed verify nonconvexity for the SINR region. However, Fig. 3(b) indicates that the rate region may still exhibit convexity. This result does not contradict Corollary 2 because the conditions we derived are sufficient but not necessary. Nevertheless, the potential convexity of the rate region, despite a nonconvex SINR region, provides numerical evidence supporting the design recommendation outlined in Remark 5.
@@
Figure 3: Sample points on weak Pareto boundary when the conditions of Corollary 2 are not satisfied.

### v3

@@ IV Discussion
To highlight further the relevance of the preceding results for network designers, except for Sect. IV-E, we focus on the interference model in Examples 6 and 8. Notably, that interference model is very common in the massive/extremely large MIMO and cell-less literature (see [26, 13, 12] and the references therein). In the discussion that follows, we extensively use Fact 4 in the Supplemental Material, which establishes that an invertible interference matrix ${\bm{M}}\in{\mathbb{R}}_{+}^{N\times N}$, being nonnegative, is an inverse Z-matrix if and only if it is an inverse M-matrix. This relationship allows us to utilize the insights from the previous sections, along with the comprehensive body of research on these matrix classes [19], to explain interference patterns observed in common network models for resource allocation.
@@ IV-A Impact of self-interference on the convexity of the feasible rate and SINR regions
The diagonal entries of the interference matrix ${\bm{M}}\in{\mathbb{R}}_{+}^{N\times N}$ defined in Example 8 correspond to the so-called self-interference, which in many models in the massive MIMO and cell-less literature is caused by the lack of perfect channel state information. For Remark 4 to be applicable, self-interference must exist in all communication links $n\in\mathcal{N}$. The reason is that a necessary (but not sufficient) condition for ${\bm{M}}\in{\mathbb{R}}_{+}^{N\times N}$ to be an inverse Z-matrix (or, equivalently, an inverse M-matrix, as mentioned above) is that all diagonal entries must be positive [19, Theorem 2.9].
@@
Fact 2 in the Supplemental Material shows that there exist scalars $(\alpha_{1},\ldots,\alpha_{K+1})\in{\mathbb{R}}_{+}^{K+1}$ such that $(\alpha_{n}{\bm{I}}+{\bm{M}}+{\bm{u}}{\bm{a}}_{n}^{t})_{n\in\{1,\ldots,K\}}$ and $\alpha_{K+1}{\bm{I}}+{\bm{M}}$ are inverse M-matrices. It then follows from Fact 3 in the Supplemental Material that, for any ${\bm{S}}\in\mathcal{D}_{+}^{N\times N}$ satisfying ${\bm{S}}\geq{\bm{S}}_{\mathrm{min}}:=\bar{\alpha}{\bm{I}}$, where $\bar{\alpha}:=\max\limits_{n\in\{1,\ldots,K+1\}}\alpha_{n}$, the matrices ${\bm{M}}+{\bm{S}}$, ${\bm{M}}+{\bm{S}}+{\bm{u}}{\bm{a}}_{1}^{t}$, …, ${\bm{M}}+{\bm{S}}+{\bm{u}}{\bm{a}}_{K}^{t}$ are inverse M-matrices, hence also inverse Z-matrices because they are nonnegative (Fact 4 in the Supplemental Material). A direct application of Remark 4, Corollary 1, and Corollary 2 completes the proof. ∎
@@ IV-B Users offering compatible channels
Assume that two single-antenna users need to be served by a single base station equipped with $L\in{\mathbb{N}}$ antennas (we consider the general case soon). In the massive MIMO literature, a key concept indicating whether both users can be served simultaneously while maintaining good link performance is favorable or approximately favorable propagation [23, Ch. 7]. More precisely, we say that the users offer favorable propagation if their channels $({\bm{h}}_{1},{\bm{h}}_{2})\in\mathbb{C}^{L}\times\mathbb{C}^{L}$ are such that ${\bm{h}}_{1}^{H}{\bm{h}}_{2}=0$, or that they offer approximately favorable propagation if $|{\bm{h}}_{1}^{H}{\bm{h}}_{2}|$ is sufficiently small (possibly in a probabilistic sense for models such as those in Example 6). It is often recommended to avoid scheduling users on the same resources when favorable propagation is absent [23, p. 154], and this principle can be trivially generalized to systems with multiple users and base stations by considering the channels for each link pair. However, this system design recommendation overlooks key aspects of real systems, including, among others: (i) it does not take into account the presence of beamformers and the available transmit power; (ii) it ignores self-interference; (iii) many results proving approximate favorable propagation consider asymptotic regimes by letting the number of antennas diverge to infinity, but systems with a small number of antennas are common; (iv) users cannot be simply dropped – they need to be scheduled at some time slot considering some fairness criterion; (v) and it lacks mathematical guarantees that users would not achieve better performance by being divided into smaller groups and served at different time slots.
The results in Sect. III provide a framework for defining a notion of user compatibility that overcomes the above drawbacks. Specifically, if the achievable rate region is convex, as discussed in Sect. I-A, we have a mathematical certificate that the network performance cannot be improved by splitting users into smaller groups and by serving these groups at different times, irrespective of the point on the weak Pareto boundary being considered (the fairness criterion). Therefore, the following definition is justified:
@@
Consider the system model in Examples 6 and 8. We say that the users offer Z-compatible channels without power constraints if the interference matrix $\bm{M}$} is an inverse Z-matrix. Likewise, we say that the users offer Z-compatible channels with the power constraint $\mathcal{P}$ in (23) if the matrices $({\bm{M}}+{\bm{ua}}_{n}^{t})_{n\in\{1,\ldots,K\}}$ are inverse Z-matrices.
@@ IV-C Numerical aspects
If the interference matrix $\bm{M}$} is of dimension $2\times 2$ (i.e., we try to produce a certificate of channel Z-compatibility for two users), inverting such a small matrix may not pose serious challenges. Nevertheless, it is more efficient and numerically stable to apply the result stated in Fact 1 in the Supplemental Material. More specifically, an interference matrix ${\bm{M}}\in{\mathbb{R}}_{+}^{2\times 2}$ is an inverse Z-matrix if and only if its determinant $\mathrm{det}({\bm{M}})$ satisfies $\mathrm{det}({\bm{M}})>0$. Unfortunately, this simple inequality is only a necessary condition for larger matrices [19, Theorem 1.2]. Nevertheless, all principal minors of inverse Z-matrices are positive [19, Corollary 2.3.3], which lead us to a simple rule for identifying pairs of users potentially leading to a nonconvex SINR or rate region:
@@
Consider the interference model in Examples 6 and 8. Let ${\bm{M}}\in{\mathbb{R}}_{+}^{N\times N}$ be the interference matrix of a network with NN users. Given two users $(i,j)\in\mathcal{N}\times\mathcal{N}$, let ${\bm{M}}[i,j]\in{\mathbb{R}}^{2\times 2}_{+}$ be the reduced interference matrix obtained from $\bm{M}$} by keeping only the $i$th and $j$th columns and the $i$th and $j$th rows. If $\mathrm{det}({\bm{M}}[i,j])\leq 0$, then the full matrix $\bm{M}$} is not an inverse Z-matrix.
IV-D A counter-example to the conjecture in [35, Sect. 5.4.4][36]
Given a matrix ${\bm{M}}\in{\mathbb{R}}_{+}^{N\times N}$, the studies in [35, Sect. 5.4.4][36] conjecture that convexity of the function $l_{\bm{M}}:{\mathbb{R}}^{N}_{+}\to{\mathbb{R}}_{+}:{\bm{x}}\mapsto\rho({\mathrm{diag}(\bm{{\bm{x}}})}{\bm{M}})$ is linked to the positive semidefiniteness of the symmetrized interference matrix ${\bm{M}}+{\bm{M}}^{t}$. If this conjecture were true, it would provide a simple sufficient condition for verifying the convexity of the SINR region $\mathcal{S}$ for the system model in Example 6. We now show that positive semidefiniteness of ${\bm{M}}+{\bm{M}}^{t}$ is neither necessary nor sufficient for convexity. More specifically, if the interference matrix ${\bm{M}}\in{\mathbb{R}}^{N\times N}_{+}$ is an inverse Z-matrix, in which case the function $l_{\bm{M}}$ is convex, then there exists a positive diagonal matrix ${\bm{D}}\in{\mathbb{R}}^{N\times N}_{+}$ such that the symmetrized interference matrix ${\bm{DM}}+{\bm{M}}^{t}{\bm{D}}$ is positive definite [19, Theorem 1.4], but $\bm{D}$} is not necessarily the identity matrix. Therefore, positive definiteness of ${\bm{M}}+{\bm{M}}^{t}$ is not a necessary condition for convexity of $l_{\bm{M}}$. To show that this property is not sufficient, take ${\bm{x}}_{1}=[0.5,~0.1,~1]^{t}$, ${\bm{x}}_{2}=[0.5,~0.5,~0.5]^{t}$, $\alpha=0.9$, and
@@ IV-E The sum-rate maximization problem
In this final subsection of Sect. IV, we turn the attention to the interference model in (20) (or, equivalently (22)) in its full generality, with the models presented in Examples 6 and 8 serving as particular cases. In more detail, many scheduling and resource allocation schemes rely heavily on the sum-rate maximization problem, which is known to be NP-hard in general [22]. However, if $T_{\|\cdot\|}$} in (25) is norm-inducing, the closure $\overline{\mathcal{R}_{\mathcal{P}}}=\{{\bm{r}}\in{\mathbb{R}}_{+}^{N}\mid\rho((e^{{\mathrm{diag}(\bm{r})}}-{\bm{I}}){T_{\|\cdot\|}})\leq 1\}$ of the set of achievable rates $\mathcal{R}_{\mathcal{P}}$ in (32) is convex as shown in Proposition 6, and $\overline{\mathcal{R}_{\mathcal{P}}}$ can be equivalently expressed using a monotone norm as shown in (34). In this case, (weighted) sum-rate maximization can be posed as the following convex optimization problem, which opens up the possibility for deriving efficient solvers that provably converge to global optima (particular solvers will be considered elsewhere):
$\begin{array}[]{rl}\mathrm{max.}&{\bm{w}}^{t}{\bm{r}}\\ \mathrm{s.t.}&{\bm{r}}\in\overline{\mathcal{R}_{\mathcal{P}}}=\{{\bm{r}}\in{\mathbb{R}}_{+}^{N}\mid\rho((e^{\mathrm{diag}(\bm{r})}-{\bm{I}}){T_{\|\cdot\|}})\leq 1\},\end{array}$ (37)
@@
It is worth emphasizing that the class of sum-rate maximization problems that can be reformulated as convex optimization problems without change in variables does not necessarily require $T_{\|\cdot\|}$} to be norm-inducing, and the results from Sect. III offer analytical support for a straightforward guideline for designing sum-rate maximization algorithms:
@@
We have shown that convexity of the function ${\bm{s}}\mapsto\rho({\mathrm{diag}(\bm{s})}{T_{\|\cdot\|}})$ implies convexity of the set $\overline{{\mathcal{S}}_{\mathcal{P}}}$ and also convexity of the set $\overline{\mathcal{R}_{\mathcal{P}}}$, as established in Proposition 6. However, $\overline{\mathcal{R}_{\mathcal{P}}}$ can be convex even if neither the set $\overline{{\mathcal{S}}_{\mathcal{P}}}$ nor the function ${\bm{s}}\mapsto\rho({\mathrm{diag}(\bm{s})}{T_{\|\cdot\|}})$ is convex. This fact suggests to use the achievable rates directly as the optimization variables in sum-rate maximization problems, as done in (37). The alternative approach of using SINR values ${\bm{s}}=(s_{1},\ldots,s_{N})$ as the optimization variables (or the power levels providing the required SINR values) and the sum-rate function $\sum_{n\in\mathcal{N}}\log(1+s_{n})$ as the cost function may require the introduction of the set $\overline{{\mathcal{S}}_{\mathcal{P}}}$ as a constraint, which is less likely to be convex compared to the set $\overline{\mathcal{R}_{\mathcal{P}}}$. Therefore, we may lose the opportunity to exploit a potential hidden convexity of the sum-rate maximization problem, increasing the risk of iterative solvers converging to local rather than global optima, if convergence can be guaranteed.
@@ V Simulations
To illustrate the results from the previous section using the model in Examples 6 and 8, we examine the uplink of a small cell-less network consisting of four access points, each equipped with two antennas, uniformly distributed within a 100m × 100m square area. Three single-antenna users are positioned uniformly at random in this area, with each user connected to the two access points providing the strongest channels. The small number of users and access points is chosen deliberately because plotting the feasible SINR and rate regions is impossible if the number of users exceeds three. Nevertheless, the results in the previous section are valid for systems of any dimension. For the beamformers, we employ maximum ratio combining. Concretely, in the SINR expression in (21), the aggregate beamformer vector ${\bm{v}}_{n}$ (obtained by stacking the beamformers of every access point) for each user $n\in\mathcal{N}=\{1,2,3\}$ is a scaled version of the aggregated channel vector ${\bm{h}}_{n}$, with zeros placed in the coordinates corresponding to the access points to which the user is not connected. Expectations are replaced by empirical averages from 100 realizations of the random variables, and all channel parameters and noise samples are constructed exactly as described in [26, Sect. III-D], so we omit the details for brevity. For the monotone norm $\|\cdot\|$ in the power constraint $\mathcal{P}$ in (23), we use the $l_{\infty}$ norm, scaled such that the maximum transmit power per user is limited to $p_{\max}=0.2$ W; i.e., we use the polyhedral monotone norm $(\forall{\bm{p}}\in{\mathbb{R}}^{N})~\|{\bm{p}}\|:=(1/p_{\max})~\|{\bm{p}}\|_{\infty}$. Equivalently, the vectors $({\bm{a}}_{n})_{n\in\{1,2,3\}}$ in Corollary 2 are the standard basis vectors appropriately scaled so that the users do not exceed the maximum transmit power.
@@
Fig. 2 illustrates results for a sample user placement scenario where all matrices $({\bm{M}}+{\bm{u}}{\bm{a}}^{t}_{n})_{n\in\{1,2,3\}}$ described in Example 6 are inverse Z-matrices. These figures indicate convexity of the feasible rate and SINR regions, which is in agreement with the results in Corollary 2.
@@
which, in light of Remark 6, cannot be an inverse Z-matrix because, in particular, the determinant of the $2\times 2$ matrix obtained by removing the second row and the second column is negative; i.e., users 1 and 3 do not offer Z-compatible channels in this three-user network. Consequently, our theoretical results cannot guarantee convexity of the SINR and rate regions, and we indeed verify nonconvexity for the SINR region. However, Fig. 3(b) indicates that the rate region may still exhibit convexity. This result does not contradict Corollary 2 because the conditions we derived are sufficient but not necessary. Nevertheless, the potential convexity of the rate region, despite a nonconvex SINR region, provides numerical evidence supporting the design recommendation outlined in Remark 7.
@@
Figure 3: Sample points on the weak Pareto boundary when the conditions of Corollary 2 are not satisfied.